# Density-Based Long-Range Electrostatic Descriptors for Machine Learning Force Fields

### v1

@@ III.3 Zirconia
The data set of this material consists of 592 distinct structures, each containing 32 Zr and 64 O atoms. The configurations are taken from VASP MD simulations. During these simulations, the material was heated between 300K to 2800K using the Langevin thermostat. For further details on the data set we refer to Ref. Verdi et al. (2021).
@@
The hyperparameters used for the short-range parts and MACE were identical to those employed for NaCl. However, for the density-based descriptor, choosing $N_{\text{max}}=5$, $s=1.5$, $\zeta_{\text{max}}=2$ yields the best results. The LODE method, when using $N_{\text{max}}=4$ and $R_{\text{cut}}=3~{}$Å instead of $N_{\text{max}}=1$ and $R_{\text{cut}}=1~{}$Å, yielded slightly better results. Therefore, these values were used in the subsequent calculations.
A comparison of the results for zirconia in Tab. 2 shows that the MACE with MP yields the smallest error. However, the purely local methods result in errors in the single-digit percentage range, indicating that the system is already well-described using only short-range descriptors. The addition of long-range entries from the density-based and LODE method does not significantly alter the results.
Table 2: The root-mean-square percentage error (RMSPE) was calculated for the validation data of ZrO2 using combined density and LODE descriptors, as well as local-only descriptors and MACE Batatia et al. (2022), with and without Message Passing (MP). For a cutoff radius of $6$ in the short-range part of the descriptor, there is hardly any difference between the three kernel methods and they do not reach the accuracy of MACE. However, the error is slightly smaller when long-range electrostatic interactions are included.
Local MACE no MP LODE Density MACE
RMSPE E (%) 1.8 1.7 1.9 1.4 1.0
RMSPE F (%) 6.8 6.0 6.3 6.2 3.3
As both the LODE and the density-based model describe only monopole-monopole interactions and yield results that are as good as those of the local description, it can be concluded that no significant monopole-monopole interactions are present in zirconia.
When comparing MACE with and without MP it is evident that the long-range MP model improves the results. We speculate that this implies that, despite the absence of monopole-monopole interactions in zirconia, higher-order long-range electrostatic interactions, such as dipole-dipole interactions, play an important role. Although our results are indirect, we suspect that MACE (as well as other message-passing networks) can describe dynamical long-range dipole-dipole interactions. Considering their architecture, they should be able to learn the dynamic, displacement-induced dipoles on one site (1st MP layer) and their interaction with other sites (2nd MP layer).
@@ IV Conclusion
MLFFs have advanced significantly over the past years, with the equivariant message passing networks in particular greatly improving prediction quality for materials where some form of long-range physics is involved. The consensus seems to be that for solid-state materials, message-passing networks improve the accuracy typically by a factor of two to three compared to more traditional invariant perceptions or standard kernel-based approaches.
@@
The exact reason for this is not yet fully understood. The present work aims to take a rational approach to the problem and attempts to propose long-range descriptors that are capable of describing a specific type of long-range interaction. Our starting point is the long-distance equivariant (LODE) approach, which is designed to describe electrostatic interactions between charges Grisafi and Ceriotti (2019) but also dipoles Grisafi, Nigam, and Ceriotti (2021); Huguenin-Dumittan et al. (2023). The important difference is that we wanted to stick to the standard kind of descriptors that express the environment of an atom in a suitable set of invariant descriptors. It turns out that this is possible and only requires one to abandon the finite-range projectors usually used to describe the environment.
@@
We show that the present long-range descriptor can achieve almost the same accuracy as the LODE descriptors for point charge models. Matter of fact, the learning efficiency is worse than for LODE, since the machine learning model has to determine from the data which linear combination of descriptors describes the Coulomb $1/r$ law. However, the present approach is more flexible as it can, by construction, describe any, e.g. a screened Coulomb, interaction. We demonstrate this for a real material, liquid NaCl, where the interactions are dominated by electrostatic interactions between point charges. However, these interactions are screened by the electrons and the screening is distance-dependent. In this case, the present density-based descriptors combined with the usual short-range descriptors outperform LODE combined with short-range descriptors. Nevertheless, the model is found to be only as accurate as MACE and cannot improve upon the flexible message-passing model.
@@
For the second test material, ZrO2, we find no improvement using long-range descriptors, and the performance of MACE cannot be matched. This indicates that we are still lacking important physics in the descriptors considered here. The likely explanation is that in ZrO2 the dynamical (Born effective) charges are strongly anisotropic, i.e. moving in the O-Zr bond direction and orthogonally to them, respectively, gives very different Born effective charges. Our surrogate model is not able to describe this by construction (it rather assumes an isotropic, possibly screened interaction).
@@
Clearly, more work is needed to fully understand what kind of physics needs to be included to improve the short-range models that have dominated research over the last decade. We believe that this rational approach remains relevant, even though data-based flexible message-passing networks are now outperforming the rationally designed surrogate models.
@@ V Acknowledgements
This research was funded in whole by the Austrian Science Fund (FWF) 10.55776/F81. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. The presented computational results have been largely obtained using the Vienna Scientific Cluster (VSC).

### v3

@@ III.4 Zirconia
The data set for this material consists of 592 distinct structures, each containing 32 Zr and 64 O atoms. The configurations are taken from VASP MD simulations. During these simulations, the zirconia was heated from 300K to 2800K using a Langevin thermostat in an NPT ensemble. For further details on the data set we refer to Ref. Verdi et al., 2021.
@@
The hyperparameters used for the short-range parts and MACE were identical to those employed for NaCl. However, for the density-based descriptor, choosing $N_{\text{max}}=5$, $s=1.5$, $\zeta_{\text{max}}=2$ yields the best results.
For the "minimal" LODE approach, again $R_{\text{cut}}=1~{}$Å  $N_{\text{max}}=1$ and $\sigma=0.3$ was used. Increasing the cutoff radius further to $R_{\text{cut}}=5~{}$Å and setting $N_{\text{max}}=5$ (LODE flexible) does not significantly improve the results. In this specific case, we also attempted to include three body descriptors with $L_{\text{max}}=1$ (LODE three-body) as anisotropic dipole interactions may play an important role (see below), but this also did not improve the results significantly, as demonstrated in the Table 3. Here for the kernel methods the short- and long-range parts of the descriptors are weighted equally, which a hyper-parameter search showed to be optimal.
A comparison of the results for zirconia in Tab. 3 shows that MACE with MP yields the smallest error. However, even the purely local methods result in errors in the single-digit percentage range, indicating that the system is already fairly well-described using only short-range descriptors. The addition of long-range entries from the density-based and LODE method does not significantly improve the results.
Table 3: The root-mean-square percentage error (RMSPE) was calculated for the validation data of ZrO2 using combined density and LODE descriptors describing the electrostatic potential, a more flexible one and a three-body one, as well as local-only descriptors and MACE Batatia et al. (2022), with and without Message Passing (MP) as well as using invariant (inv.) and equivariant (equiv.) features. For a cutoff radius of $6$ in the short-range part of the descriptor, there is hardly any difference between the three kernel methods and they do not reach the accuracy of MACE.
RMSPE E (%) RMSPE F (%)
Local 1.8 6.8
MACE no MP 1.7 6.0
LODE minimal 1.7 6.5
LODE flexible 1.8 6.2
LODE three-body 1.9 6.3
Density 1.4 6.2
MACE inv. 1.0 3.3
MACE equiv. 0.7 2.4
As both the LODE (with two-body descriptors) and the density-based model describe only monopole-monopole interactions and yield results that are as good as those of the local description, it can be concluded that no significant monopole-monopole interactions are present in zirconia.
When comparing MACE with and without MP it is evident that the MP model improves the results. We speculate that, despite the absence of simple monopole-monopole interactions in zirconia, higher-order long-range electrostatic interactions, such as dipole-dipole interactions, play an important role. Although our results are indirect, we suspect that MACE (and other message-passing networks) can describe dynamical long-range dipole-dipole interactions. Specifically, MP networks can learn dynamic, displacement-induced and direction-dependent dipoles on one site in early layers (e.g. MACE 1st MP layer) and deduce the interaction energies with other sites in later layers. We cannot see how a simple density-based descriptor or LODE can potentially model such interactions, and our empirical evidence also points against it.
@@ IV Conclusion
MLFFs have advanced significantly over the past years, with the equivariant message passing networks in particular greatly improving prediction quality for materials where some form of long-range physics is involved. The consensus seems to be that for solid-state materials, message-passing networks improve the accuracy typically by a factor of two to three compared to more traditional invariant perceptrons or standard kernel-based approaches.
@@
The exact reason for this is not yet fully understood. The present work aims to take a rational approach to the problem and attempts to propose long-range descriptors that are capable of describing a specific type of long-range interaction. Our starting point is the long-distance equivariant (LODE) approach, which is designed to describe electrostatic interactions between charges Grisafi and Ceriotti (2019) but also static dipoles Grisafi, Nigam, and Ceriotti (2021); Huguenin-Dumittan et al. (2023). The important difference is that we aimed to stay within the standard scheme to express the environment of each atom in a suitable set of invariant descriptors. It turns out that this is possible and only requires one to adopt suitable slowly decaying projectors.
@@
We show that the present long-range descriptor can achieve almost the same accuracy as the LODE descriptors for point charge models. Matter of fact, the learning efficiency is worse than for LODE, since the machine learning model has to determine from the data which linear combination of descriptors describes the Coulomb $1/r$ behavior. However, the present approach is more flexible as it can, by construction, describe any interaction, e.g. even a distance-dependent screened Coulomb interaction. We demonstrate this for a real material, liquid NaCl, where the long-range interactions are dominated by electrostatic screened interactions between point charges. For this particular test case, the present density-based descriptors combined with the usual short-range descriptors appear to outperform LODE combined with short-range descriptors. Nevertheless, the model is found to be only as accurate as MACE and cannot improve upon the flexible message-passing model. For solid NaCl, LODE and the present long-range density descriptors perform very similarly and MACE clearly outperforms them.
@@
For the second test material, various phases of ZrO2, we find no improvement using long-range descriptors, and the performance of MACE cannot be matched. This indicates that the present electrostatic models (long-ranged density approach as well as LODE) still lack important physics. The likely explanation is that in ZrO2 the dynamical (Born effective) charges are strongly anisotropic, i.e. moving in the O-Zr bond direction or orthogonally to them, respectively, gives very different Born effective charges. Such a behavior cannot be approximated using monopole-monopole interactions only, and hence our surrogate model is unable to describe this by construction (it rather assumes an isotropic, possibly distance dependent interaction).
@@
Clearly, more work is needed to fully understand what kind of physics needs to be included to improve the short-range models that have dominated research over the last decade. We believe that this rational approach remains relevant, even though data-based flexible message-passing networks are currently outperforming the rationally designed surrogate models.
@@ V Acknowledgements
This research was funded in whole by the Austrian Science Fund (FWF) 10.55776/F81. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. The presented computational results have been largely obtained using the Vienna Scientific Cluster (VSC).
VI Author Declarations
VI.1 Conflict of Interest
The authors have no conflicts to disclose.
VII Data Availability
The datasets used/produced in this work can be freely accessed at https://doi.org/10.5281/zenodo.13985979.